OpenStudy (anonymous):

Hockey teams receive 2 points when they win and 1 point when they tie. One season, a team won a championship with 48 points. They won 9 more games than they tied. How many wins and ties did they have?

OpenStudy (anonymous):

you need to get two equations out of this question to solve

OpenStudy (anonymous):

we have two key pieces of information that tell us what these two equation will look like: they recieved 48 points at the end of the season, they won 9 more games than they tied

OpenStudy (anonymous):

let us use the variables W for win and T for tie

OpenStudy (anonymous):

our first equation will be derrived from the 48 points piece of data. since wins are worth 2 points and ties are worth 1, our first equation will be 2*W+T=48

OpenStudy (anonymous):

next, we know that they won 9 more games than they tied, so we also know W=T+9 or, to make it easier to plug in, T=W-9

OpenStudy (anonymous):

if you plug this second equation into the first one, what do you get?

OpenStudy (anonymous):

?

OpenStudy (anonymous):

we have two equations, 1) 2W+T=48 2) T=W-9 we are going to substitute the second equation into the first for T

OpenStudy (anonymous):

look at it like this 2*W+(T)=48

OpenStudy (anonymous):

now substituting 2W+(W-9)=48

OpenStudy (anonymous):

all i did was replace T with (W-9)

OpenStudy (anonymous):

now we remove the brackets 2W+W-9=48

OpenStudy (anonymous):

combine the W terms 3W-9=48

OpenStudy (anonymous):

add 9 to both sides 3W=57

OpenStudy (anonymous):

and divide by 3 W=19

OpenStudy (anonymous):

voila, we now know how many games they won in the season, 19

OpenStudy (anonymous):

now, remembering the identity we used for T T=W-9 substitute this value in for W

OpenStudy (anonymous):

10

OpenStudy (anonymous):

you got it

OpenStudy (anonymous):

for any problem like this, you will have these same steps to follow

OpenStudy (anonymous):

1) get two equations 2) use one of these equations to get an identity for one variable in terms of the other 3) substitute the identity into the the other equation 4) solve thats all there is to it